The locus of pole of chord of parabola which subtend right angle  vertex 

Grade:11

1 Answers

Arun
6 years ago
Dear Prajwal
 
ler the pole be (x1,y1)  then polar is yy1-2a(x+x1) on homogenising with parabola we get  homogenous pair of straight lines ie           8a2-4ay1xy+2ax1y2=0 since it subtend right angle at vertex then coefficient of x2+ coefficient of y2 =0 therefore 8a2+2ax1 =0 therefore locus is x1+4a=0 , x+4a=0
 
 
Regards
Arun (askIITians forum expert)